Lead nitrate can be decomposed by heating. What is the percent yield of the decomposition
reaction if 10.9 g Pb(NO ) are heated to give 5.9 g of PbO?
2Pb(NO ) (s) 2PbO(s) + 4NO (g) + O (g)
1
Expert's answer
2016-03-09T08:36:42-0500
The theoretical mass of PbO:
m(PbO) = ((2*M(PbO))*m(2Pb(NO3)2))/(2*M(2Pb(NO3)2)) = (2*223 g/mol*10.9 g)/(2*331 g/mol) = 7.3 g
The percent yield of the decomposition reaction:
X = (5.9 g/7.3 g)*100% = 80.8 %
Answer:
The percent yield of the decomposition reaction:
80.8 %